OpenStudy (anonymous):

$6,000 for 6 years at 8½% compounded daily will grow to?

OpenStudy (anonymous):

whatever \(6000(1+\frac{.085}{365})^{6\times 365}\) is need a calculator for sure

OpenStudy (anonymous):

ok let me see im going to try it on mathematics microsoft

OpenStudy (anonymous):

What is the little 6x365? I know that the 6000(1+.085/365) = 6001.40

OpenStudy (istim):

Compound interest formula? If so, just plug and play.

OpenStudy (anonymous):

don't round, put it in your calculator all at once the only thing you should compute first is \(6\times 365=2190\)

OpenStudy (anonymous):

you should compute \(6000\times (1+\frac{.085}{365})^{2190}\) in one step you are taking such a large power, that any little truncation makes a huge difference at the end

OpenStudy (anonymous):

the answer was $3628.20

OpenStudy (anonymous):

that is not possible since you started with $6,000